OpenStudy (georgia412):

Find the equation for the straight lines which passes through points (-7,3) and (2,4).. I think x= 1/9 and y might = 1 1/9?

OpenStudy (zzr0ck3r):

\[slope = \frac{x_1-x_0}{y_1-y_0}\] after you find slope, do the same thing we did last time ( you can choose either point) to find b

OpenStudy (georgia412):

That's what I did.. Is this the right answer??

OpenStudy (georgia412):

Which the equation would be y=1/9x+1 1/9?

OpenStudy (georgia412):

*1 3/9 sorry

OpenStudy (zzr0ck3r):

ahh one sec m = 1/9 y = (1/9)x+b so 4=(1/9)2+b 4=(2/9)+b b=4-(2/9) b=36/9-2/9 b=34/9 y = (1/9)x+(34/9)

OpenStudy (zzr0ck3r):

you can also use (y-y_0)=(x-x_0)m y-4=(x-2)(1/9) y = (1/9)x-(2/9)+4 y=(1/9)x-(2/9)+36/9 y = (1/9)x+(34/9)

OpenStudy (georgia412):

Thanks! I just left out the last step :)